OpenStudy (anonymous):

Need help finding the quartic function with only two real given x=-1 and x=-3

OpenStudy (anonymous):

(x+1)(x+3) = x^2 + 3x + x + 3 x^2 + 4x + 3

OpenStudy (anonymous):

how do I find the answer

OpenStudy (anonymous):

isnt that the answer that I gave

OpenStudy (anonymous):

doesnt match any that they give

OpenStudy (anonymous):

is it quadratic or quartic?

OpenStudy (anonymous):

y=x^4-4x^3-4x^2-4x-3 y=-x^4+4x^3+4x^2+4x+3 y=x^4+4x^3+3x^2+4x-4

OpenStudy (anonymous):

quartic function

OpenStudy (anonymous):

do you know how to do synthetic division?

OpenStudy (anonymous):

no we are learning but I dont understand it

OpenStudy (anonymous):

its easy! check one by one: first x=-1 y=x^4-4x^3-4x^2-4x-3 -1) 1 -4 -4 -4 -3 -1 5 -1 5 -------------------- 1 -5 1 -5 2

OpenStudy (anonymous):

would it be y=x^4+4x^3+3x^2+4x-4

OpenStudy (anonymous):

so y=x^4-4x^3-4x^2-4x-3 -----> u can eliminate this

OpenStudy (anonymous):

lets check

OpenStudy (anonymous):

i am sorry it si hard to do division on here!

OpenStudy (anonymous):

can you explain how you got the numbers in the first one you did

OpenStudy (anonymous):

ok first you gave me two real zeros: x= -1 and x= -3 I took the first zero x = -1

OpenStudy (anonymous):

and then put it in divisor's place

OpenStudy (anonymous):

then, the numbers after the parentheses are the coefficients of the terms of one of the choices you gave me, following me?

OpenStudy (anonymous):

i think so for like the one i picked it would be -1[1 4 3 4 4 but not sure what the numbers on bottom would be

OpenStudy (anonymous):

the last 4 would be negative cuz its in the equation

OpenStudy (anonymous):

-1) 1 4 3 4 -4

OpenStudy (anonymous):

Yea I seen I made that mistake after I sent it

OpenStudy (anonymous):

yeah there, so far you good?

OpenStudy (anonymous):

but how did you get the numbers on bottom

OpenStudy (anonymous):

ok i will tell you but before that let me tell you one more thing when lining up the coefficients.....for the equation we picked it has all the terms from degree 4, not missing a degree in the middle! But what if its like x^4 + x^2 + 3? what would you do then?

OpenStudy (anonymous):

would they both be 1 1 3

OpenStudy (anonymous):

no! zero would take place in the missing places 1 0 1 0 3

OpenStudy (anonymous):

okay I see I was putting 1s for the xs

OpenStudy (anonymous):

no no, I meant x^4 + x^2 + 3 it would be x^4 + 0x^3 + x^2 + 0x + 3 I was talking about those zeros

OpenStudy (anonymous):

yea i see that i knew what you meant after i put mine cause you have to have 5 numbers

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now those very bottom numbers: -1) 1 4 3 4 -4 0 --------------------- First you bring down the first coefficient down, imaging that theres a zero under the first 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then add 1 + 0 -1) 1 4 3 4 -4 0 ----------------------- 1 then multiple the bottom 1 with the divisor, which in this case is -1 1(-1) = -1 -1) 1 4 3 4 -4 0 -1 ---------------------------- 1 the product goes int he middle line after the 0

OpenStudy (anonymous):

am i confusing you?

OpenStudy (anonymous):

not yet Im following

OpenStudy (anonymous):

ok, then you would the same with 4 and -1 as you did with 1 and 0: add 4 and -1

OpenStudy (anonymous):

3

OpenStudy (anonymous):

the answer will take place in the bottom line after 1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then you multiple 3 with the divisor, which is -1 and youll get -3 -3 will take place under the coefficient 3 in the middle line

OpenStudy (anonymous):

want me to go on?

OpenStudy (anonymous):

stop me whn u get confused, kay?

OpenStudy (anonymous):

if you would please

OpenStudy (anonymous):

so under the -4 it would be 4

OpenStudy (anonymous):

then add 3 and -3 we will get 0, which will take its place in the bottom line

OpenStudy (anonymous):

so far -1) 1 4 3 4 -4 0 -1 -3 -1 --------------------------------- 1 3 0

OpenStudy (anonymous):

good?

OpenStudy (anonymous):

yea I had a wrong number on the top line

OpenStudy (anonymous):

ok then you add 4 and -1, which is 3 this 3 will take place in the bottom line after 0

OpenStudy (anonymous):

then we multiply the 3 with the divisor, -1 and the answer is -3

OpenStudy (anonymous):

ok now im lost

OpenStudy (anonymous):

-1) 1 4 3 4 -4 0 -1 -3 -1 ------------------------- 1 3 0 3

OpenStudy (anonymous):

wait a sec! brb!

OpenStudy (anonymous):

how come you dont add -4 to -1 and add to middle row

OpenStudy (anonymous):

are u talking about the coeffiecient -4?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

we havent gotten there yet!!

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

-1) 1 4 3 4 -4 0 -1 -3 -1 -3 ---------------------------- 1 3 0 3

OpenStudy (anonymous):

following?

OpenStudy (anonymous):

how did you get -3

OpenStudy (anonymous):

whats the last number on the bottom line?

OpenStudy (anonymous):

it is 3 right? so I multiplied it with -1 and I got -3 which put in the middle line under -4

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

then you add -4 and -3, and we get -7

OpenStudy (anonymous):

whenever we get a number instead of zero in the bottom, we got -7, so its not a zero for that equation

OpenStudy (anonymous):

doing ok?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

here we got -7, but the question says that the function has to include x = -1 right? suppose this x = -1 is the zero/root/xint, we would have gotten zero instead of -7

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

we can take that option out

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do the same with the other 2 equations and let me know

OpenStudy (anonymous):

ok I will

OpenStudy (anonymous):

i am going to look around, call me after u r done! :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

not sure if Im going in the right direction -1 1 4 3 4 -4 -2 3 -3

OpenStudy (anonymous):

hey! we already did that!!

OpenStudy (anonymous):

check the other 2 equations!

OpenStudy (anonymous):

I thought we did the one that was y=x^4+4x^3+3x^2+4x-4 and I was doing y=-x^4+4x^3+4x^2+4x+3

OpenStudy (anonymous):

I will finish tomorrow gotta get up early

OpenStudy (anonymous):

want me to tell u the answer

OpenStudy (anonymous):

y=-x^4+4x^3+4x^2+4x+3

OpenStudy (anonymous):

thats the answer!! :)

OpenStudy (anonymous):

ok!! Sorry if kept u late!! I am also studying the same

OpenStudy (anonymous):

oh, you will get it, just practice more!! u will get it for sure!! :) :) :)

OpenStudy (anonymous):

nope, but I learnt all that by myself when I was homeschooled 2 years ago

OpenStudy (anonymous):

my home teacher was an airhead

OpenStudy (anonymous):

very stupid that I taught him

OpenStudy (anonymous):

awww.....hope ur family gets better soon!! but did you go to school before? what grade are you in btw?

OpenStudy (anonymous):

same situation here.....everyone in my family caught virus including me

OpenStudy (anonymous):

oh ok!! Call me whenever you need help with other subjects!! I will see what I can do!!! started college this year!! graduated from highschool in 2011!!

OpenStudy (anonymous):

we got many things in common, I was home schooled when I was in 11th grade!! Coool!!

OpenStudy (anonymous):

oh raelly? its really nice of you to help 'em!! I was sick at the time so i had to be home schooled

OpenStudy (anonymous):

no problem! :) night!! take care!!! :))